OpenStudy (anonymous):

Solve by completing the square. Give irrational roots in simplets radical form. y^2-24y+23=0

OpenStudy (lone001):

y^2-24y+23=(y-23)(y-1)=0 y-23=0 y_{1}=23 y-1=0 y_{2}=1

OpenStudy (lone001):

y^2-24y+23=(y-23)(y-1)=0 y-23=0 y1=23 y-1=0 y2=1